# Not All Tokens Are Meant to Be Forgotten

### v1

@@ Abstract
Large Language Models (LLMs), pre-trained on massive text corpora, exhibit remarkable human-level language understanding, reasoning, and decision-making abilities. However, they tend to memorize unwanted information, such as private or copyrighted content, raising significant privacy and legal concerns. Unlearning has emerged as a promising solution, but existing methods face a significant challenge of over-forgetting. This issue arises because they indiscriminately suppress the generation of all the tokens in forget samples, leading to a substantial loss of model utility. To overcome this challenge, we introduce the Targeted Information Forgetting (TIF) framework, which consists of (1) a flexible targeted information identifier designed to differentiate between unwanted words (UW) and general words (GW) in the forget samples, and (2) a novel Targeted Preference Optimization approach that leverages Logit Preference Loss to unlearn unwanted information associated with UW and Preservation Loss to retain general information in GW, effectively improving the unlearning process while mitigating utility degradation. Extensive experiments on the TOFU and MUSE benchmarks demonstrate that the proposed TIF framework enhances unlearning effectiveness while preserving model utility and achieving state-of-the-art results. Our code is available at: https://github.com/xzhou98/LLM-Unlearning-TPO
Figure 1: Comparison of our TPO and NPO zhang2024negative on key metrics: forget quality and model utility. The right panel illustrates example responses generated by the unlearned model with TPO (ours) and NPO on the retain set. The results are derived from the Forget05 task of the TOFU dataset maini2024tofu , with the models trained for 10 epochs.
1 Introduction
Large Language Models (LLMs), pre-trained on vast text corpora, demonstrate exceptional capabilities in text generation and nuanced comprehension of language tasks brown2020language . However, LLMs exhibit a tendency to memorize parts of their training data carlini2021extracting . While memorization can be beneficial for tasks like question answering brown2020language and code generation jiang2024survey , it also raises significant security and safety concerns. Specifically, training data that includes personally identifiable information (PII) or copyrighted content poses risks of privacy violations or copyright infringement carlini2021extracting ; karamolegkou2023copyright ; li2024wmdp . To mitigate these issues, research efforts have focused on developing machine unlearning techniques to enable LLM dememorization.
@@ 1 Introduction
Machine unlearning cao2015towards ; bourtoule2021machine ; nguyen2022survey was developed as an efficient approach to remove the influence of specific training samples from (pre-)trained models, eliminating the need for full retraining. Recently, unlearning techniques have been adapted for LLMs, raising ongoing challenges in precisely removing private or copyrighted content learned from specific training samples  yao2023large ; pawelczyk2023context ; eldan2023s ; chen2023unlearn . Early approaches for unlearning LLMs focused on fine-tuning the parameters of pre-trained models li2024wmdp ; chen2023unlearn ; si2023knowledge ; jang2022knowledge . Some methods, such as the one proposed in yao2023large , often rely on gradient ascent (GA) based optimization to achieve the goal of unlearning. Since the loss function generally has no upper bound, this approach makes the precise control over parameter updates difficult, frequently resulting in catastrophic collapse, where the overall performance of the unlearned model deteriorates substantially zhang2024negative . To address this issue, methods like Negative Preference Optimization (NPO) zhang2024negative and SimNPO fan2024simplicity have developed preference optimization-based frameworks that leverage negative examples to mitigate performance collapse.
@@
Despite these foundational efforts, LLM unlearning still faces several critical challenges: (C1) Ambiguous Unlearning Targets. Most existing approaches treat the entire forget sample as the unlearning target without differentiating between unwanted information (to be unlearned) and general information (to be retained), as shown in Figure 2. This lack of distinction often results in significant degradation of model utility wang2024selective ; liu2024large ; lynch2024eight ; rezaei2024restor . (C2) Lack of Flexible and Generalizable Unwanted Information Identification. Recent methods attempt fine-grained unlearning but face critical limitations in information identification: ECO liu2024large employs a sentence-level identifier that overfits to specific keywords in the forget sample (e.g., the “college”), rather than aligning with the unlearning requester’s intent thaker2024position . This leads to insufficient unlearning when prompts containing these specific keywords are removed. While SEUL wang2024selective improves precision by leveraging generative models (e.g., ChatGPT) to identify continuous sensitive spans (e.g., PII), it remains limited to handle discontinuous and/or diverse unlearning targets (e.g., copyrighted content). This rigidity in identification compromises both effectiveness and generalizability. (C3) Sensitivity to Forget Set Size. Methods based on preference optimization zhang2024negative ; rafailov2024direct mitigate catastrophic collapse more effectively than other baselines, helping to preserve model utility. However, their effectiveness declines significantly as the forget set size increases, resulting in notable utility loss liu2024large , as shown in Figure 1.
@@
To overcome the challenges (C1)-(C3), we introduce the Targeted Information Forgetting (TIF) framework, as illustrated in Figure 2. The key contributions of our work are listed below:
@@
(1) TIF Framework. To tackle (C1), we propose a novel TIF framework for LLM unlearning. Different from existing unlearning approaches such as NPO, which predominantly unlearn entire information associated with the forget instances (e.g., $\mathbb{A_{f}}$ in the left panel of Figure 2), our TIF is designed to unlearn only the targeted unwanted information, such as the city of born in the right panel’s example. General information is often associated with some “General Words (GW)”, including stop words and commonly used phrases, which frequently appear in both retain and forget sets. In contrast, “Unwanted Words (UW)” correspond to specific private or copyrighted content, such as city of born. By specifically targeting only UW for unlearning, our TIF preserves more general information compared to existing methods like NPO, effectively preventing over-forgetting and enabling the model to generate more complete and accurate responses, as demonstrated in the retain set answers in Figure 2.
@@
(2) Unwanted Information Identification. To address (C2), we develop flexible yet effective approaches for unwanted information identification: a generative model such as ChatGPT-4, and a discriminative model such as DistilBERT sanh2019distilbert , to effectively differentiate UW from GW. We evaluate the unlearning performance of both approaches and demonstrate the use cases for each. As a bottom line, even identifying function words (e.g., the, is, or an) as GW according to linguistics would improve model utility preservation.
@@
(3) A Novel Optimization Method to Retain Model Utility. To overcome (C3), we advance preference optimization algorithms  zhang2024negative ; fan2024simplicity by introducing Targeted Preference Optimization (TPO), a novel optimization objective designed to mitigate the significant utility degradation observed in NPO. Specifically, our TPO integrates two innovative components: Preservation loss (PL) to maintain general model utility by retraining on GW, and Logit preference loss (LPL) to unlearn unwanted information in UW. This optimization approach effectively balances general information retention and unwanted information forgetting, improving the robustness of preference-guided optimization even with larger forget sets. As shown in the left panel of Figure 1, our approach, TPO, achieves a comparable forget quality to NPO while significantly preserving a higher model utility. This allows the model to generate accurate information for answers in the retain set. In contrast, NPO struggles to retain essential knowledge from the retain set, as illustrated in the right panel of Figure 1.
2 Problem Formulation
2.1 LLM Unlearning
@@
To achieve this goal, the unlearning procedure incorporates a forgetting objective on $\mathcal{D}_{f}$ and a retention objective on $\mathcal{D}_{r}$. Formally, the unlearning process is defined as yao2023large ; fan2024simplicity :
$\textstyle\min_{\boldsymbol{\theta}}\mathbb{E}_{\xi_{f}\sim\mathcal{D}_{f}}[\ell_{f}(y_{f}|x_{f};\boldsymbol{\theta})]+\mathbb{E}_{\xi_{r}\sim\mathcal{D}_{r}}[\ell_{r}(y_{r}|x_{r};\boldsymbol{\theta})],$ (1)
where $\ell_{f}$ and $\ell_{r}$ represent the forget and retain losses, respectively. Specifically, the forget loss $\ell_{f}$ determines how well the model $\mathcal{M}_{\theta}$ suppresses the association between $x_{f}$ and $y_{f}$, ensuring unwanted information is unlearned. Meanwhile, the retain loss $\ell_{r}$ enhances the model’s ability to maintain accurate associations between $x_{r}$ and $y_{r}$, preserving its original performance on $\mathcal{D}_{r}$.
2.2 Targeted Unlearning
As discussed earlier, a majority of works si2023knowledge ; chen2023unlearn ; zhang2024negative ; fan2024simplicity have consistently treated the entire token sequence $y_{f}$ as the unlearning target for each sample $\xi_{f}\coloneqq(x_{f},y_{f})$ in the forget set $\mathcal{D}_{f}$, overlooking a critical question central to the process of LLM unlearning. {mdframed}[backgroundcolor=gray!10, linecolor=black]
Critical Question for LLM Unlearning: Are all the words in the forget sample essential for unlearning in LLMs?
@@ 2.2 Targeted Unlearning
$\textstyle\min_{\boldsymbol{\theta}}\mathbb{E}_{\xi_{f}\sim\mathcal{D}_{f}}[\ell_{f}(\hat{y}|x_{f};\boldsymbol{\theta})]+\mathbb{E}_{\xi_{r}\sim\mathcal{D}_{r}}[\ell_{r}(y_{r}|x_{r};\boldsymbol{\theta})],$ (2)
@@
3 Targeted Information Forgetting (TIF)
@@
3.1 Unwanted Information Identification
@@
3.1.1 Discriminative Encoder-Only LM.
[…]
3.1.2 Generative Decoder-Only LM.
To harness the power of generative decoder-only LMs in capturing contextual and semantic information from text, we employ ChatGPT-4 to directly distinguish between UW and GW by analyzing the semantics of $y_{f}$, as shown in the right panel of Figure 3. Detailed task instructions can be found in Table 6 in the Appendix E. Furthermore, we also present a detailed comparison of discriminative and generative approaches in relation to unlearning performance in Appendix C.
3.2 Targeted Preference Optimization (TPO)
Motivation. Although numerous unlearning methods, such as NPO zhang2024negative , have demonstrated strong performance on benchmarks such as TOFU maini2024tofu and MUSE shi2024muse , most of them struggle with model utility degradation when handling large forget sets liu2024large . Specifically, we evaluate the model utility and forgetting quality of NPO on TOFU, as illustrated in Figure 5. A significant decline in model utility is evident, with the score dropping from 0.76 to 0.11, highlighting a severe utility degradation issue. We hypothesize that this degradation stems from NPO’s indiscriminate handling of the entire forget samples as unlearning targets, failing to differentiate between unwanted and general information.
@@ 3.2 Targeted Preference Optimization (TPO)
Figure 4: Model utility vs forget quality across various methods on the TOFU forget05 task. Each line represents evaluations conducted at every epoch (6 steps). “–GPT” denotes the use of ChatGPT-4 for unwanted information identification, while “PL” refers to the approach plus the Preservation Loss.
Figure 5: (a) Evaluation of loss values on GW for NPO-GPT and NPO-GPT+PL. The cross-entropy loss values for GW are plotted at each unlearning step. (b) & (c) The KL divergence between the reference model and the unlearned models on both forget and retain sets. All results are obtained for the Forget05 task in the TOFU dataset, with models trained over 10 epochs.
Preservation Loss (PL). To further investigate the reason for the model utility degradation observed in NPO-GPT, we analyze the cross-entropy loss values for GW, while the unlearning is limited to UW, as defined in Equation (2). Although the model is not explicitly optimized to forget GW, the increase in loss values for NPO-GPT indicates that GW are also forgotten, as shown in Figure 5(a). This observation naturally leads to our key idea: how can we prevent the forgetting effects on GW and, in turn, keep their loss values as low as possible during unlearning optimization? To address this, we introduce the Preservation Loss (PL), which integrates a cross-entropy loss term on GW to explicitly prevent the model from forgetting general information, formally:
@@
To validate the effectiveness of the PL term, we integrate it into NPO-GPT, forming NPO+PL-GPT, and evaluate its performance. As illustrated in Figure 5(a), incorporating PL helps maintain stable and low loss values for GW. Consequently, NPO+PL-GPT exhibits a significantly slower decline in model utility while achieving comparable forget quality to NPO and NPO-GPT, as shown in Figure 5. These initial results demonstrate that PL effectively mitigates model utility degradation, particularly for preserving the general information we aim to retain.
Logit Preference Loss (LPL). Upon further examination of Figure 5, we observe that model utility degradation persists even after incorporating unwanted information identification (GPT) and PL into NPO, as seen in NPO+PL-GPT. We hypothesize that this issue stems from the limitations of NPO itself. Specifically, the unlearning process in NPO likely introduces excessive changes from the reference model $\mathcal{M}_{\text{\bf ref}}$ (the original model before unlearning) to the final unlearned model $\mathcal{M}_{\boldsymbol{\theta}}$, particularly affecting certain general information.
To further validate this hypothesis, we analyze the logit distribution by computing the KL divergence between $\mathcal{M}_{\text{\bf ref}}$ and $\mathcal{M}_{\boldsymbol{\theta}}$ for both the forget and retain sets, as shown in Figure 5(b), (c). The high KL divergence observed in both sets suggests that $\mathcal{M}_{\boldsymbol{\theta}}$ (NPO-GPT+PL) has unintentionally forgotten not only the unwanted information but also general information. Specifically, NPO functions by directly reducing the probability assigned to target tokens, which is computed using the softmax function: $P(y_{t})=\frac{\exp{(z_{t})}}{\sum^{V}_{j=1}\exp(z_{j})}$, where the $V$ represents the vocabulary size, and $z_{t}$ denotes the logit for target token $y_{t}$. However, reducing the probability of target token $P(y_{t})$ can be achieved not only by decreasing its logits $z_{t}$ but also by increasing the logits of other tokens in the vocabulary. This unintended effect distorts the model’s overall logit distribution, potentially compromising its ability to retain general information. The key challenge is to develop a new optimization strategy for the target tokens (UW) that selectively impacts their logit distribution while preserving the general information from GW.
@@
$\ell_{\text{LPL}}(\boldsymbol{\theta})=-\mathbb{E}_{\xi_{f}\sim D_{f}}\bigg{[}\frac{2}{\beta}\log\sigma\bigg{(}\beta\frac{1}{|\hat{y}|}\sum_{i=1}^{|\hat{y}|}(z_{i}^{\text{\bf ref}}-z_{i}^{\boldsymbol{\theta}})\bigg{)}\bigg{]}.$ (4)
$z_{i}$ here denotes the logit of target token $\hat{y}_{i}$, and $\boldsymbol{\theta}$ and ref represent the parameters of the unlearned model $\mathcal{M}_{\boldsymbol{\theta}}$ and the reference model $\mathcal{M}_{\text{\bf ref}}$, respectively.
Different from NPO, LPL explicitly reduces only the logits of target tokens (UW) by enforcing a preference loss between $\mathcal{M}_{\boldsymbol{\theta}}$ and $\mathcal{M}_{\text{\bf ref}}$. The primary function of LPL is to maximize the difference in target token logits between $\mathcal{M}_{\boldsymbol{\theta}}$ and $\mathcal{M}_{\text{\bf ref}}$, while preserving the overall logit distribution. This targeted approach ensures that only the unwanted information associated with the target tokens is unlearned, without affecting general information across other tokens. As a result, LPL enables a more precise unlearning process while significantly improving model utility retention.
@@
$\mathbb{E}_{\xi_{f}\sim D_{f}}\bigg{[}\!\underbrace{-\frac{2}{\beta}\log\sigma\big{(}\beta(z_{\text{\bf ref}}(\hat{y}|x_{f})-z_{\boldsymbol{\theta}}(\hat{y}|x_{f})\big{)}}_{\text{LPL}}-\underbrace{\log P_{\boldsymbol{\theta}}(\bar{y}|x_{f})}_{\text{PL}}\!\bigg{]},$ (5)
@@
Finally, as the initial results shown in Figure 5(b), (c), our approach TPO-GPT minimizes the disruption to the logit distribution on both the forget and retain sets compared to NPO+PL-GPT. Furthermore, TPO-GPT maintains the most of model utility while achieving a comparable level of forget quality to the NPO-based methods, as shown in Figure 5.
4 Experimental Setting
4.1 Datasets and Metrics
@@
We evaluate the proposed approach alongside the baseline methods on the two widely used benchmark datasets: MUSE shi2024muse and TOFU maini2024tofu .

### v2

@@ Abstract
Large Language Models (LLMs), pre-trained on massive text corpora, exhibit remarkable human-level language understanding, reasoning, and decision-making abilities. However, they tend to memorize unwanted information, such as private or copyrighted content, raising significant privacy and legal concerns. Unlearning has emerged as a promising solution, but existing methods face a significant challenge of over-forgetting. This issue arises because they indiscriminately suppress the generation of all the tokens in forget samples, leading to a substantial loss of model utility. To overcome this challenge, we introduce the Targeted Information Forgetting (TIF) framework, which consists of (1) a flexible targeted information identifier designed to differentiate between unwanted words (UW) and general words (GW) in the forget samples, and (2) a novel Targeted Preference Optimization approach that leverages Logit Preference Loss to unlearn unwanted information associated with UW and Preservation Loss to retain general information in GW, effectively improving the unlearning process while mitigating utility degradation. Extensive experiments on the TOFU and MUSE benchmarks demonstrate that the proposed TIF framework enhances unlearning effectiveness while preserving model utility and achieving state-of-the-art results.
Codehttps://github.com/xzhou98/Unlearning-TPO
Introduction
Large Language Models (LLMs), pre-trained on vast text corpora, demonstrate strong capabilities in text generation and nuanced language understanding (brown2020language). However, they often memorize parts of their training data (carlini2021extracting), which, while useful for tasks such as question answering (brown2020language) and code generation (jiang2024survey), raises security and safety concerns. Specifically, memorization of personally identifiable information (PII) or copyrighted content poses risks of privacy violations or copyright infringement (carlini2021extracting; karamolegkou2023copyright; li2024wmdp; zhou2023hijacking; zhou2024learning; zade2025automatic). To address these risks, recent work has explored machine unlearning techniques for LLM dememorization.
@@ Introduction
Machine unlearning (cao2015towards; bourtoule2021machine; nguyen2022survey) was developed as an efficient approach to remove the influence of specific training samples from pre-trained models, eliminating the need for full retraining. Recently, unlearning techniques have been adapted for LLMs, raising ongoing challenges in precisely removing private or copyrighted content learned from specific training samples  (yao2023large; pawelczyk2023context; eldan2023s; chen2023unlearn). Early approaches mainly rely on fine-tuning pre-trained models (li2024wmdp; chen2023unlearn; si2023knowledge; jang2022knowledge), with some using gradient ascent (GA) optimization to achieve the goal of unlearning (yao2023large). However, due to the unbounded nature of the loss function, GA-based methods lack precise control over updates, often leading to catastrophic collapse, where the model’s overall performance deteriorates (zhang2024negative). To address this issue, recent work introduces preference optimization-based frameworks such as Negative Preference Optimization (NPO) (zhang2024negative) and SimNPO (fan2024simplicity) to mitigate performance collapse.
@@
Despite these foundational efforts, LLM unlearning still faces several critical challenges: (C1) Ambiguous Unlearning Targets. Most existing approaches treat the entire forget sample as the unlearning target without differentiating between unwanted information (to be unlearned) and general information (to be retained), as shown in Figure 1. This lack of distinction often leads to significant degradation of model utility (wang2024selective; liu2024large; lynch2024eight; rezaei2024restor). (C2) Lack of Flexible and Generalizable Unwanted Information Identification. Recent methods attempt fine-grained unlearning but face critical limitations in information identification: ECO (liu2024large) employs a sentence-level identifier that overfits to specific keywords in the forget sample (e.g., the “college”), rather than aligning with the unlearning requester’s intent (thaker2024position). This leads to insufficient unlearning when prompts containing these specific keywords are removed. While SEUL (wang2024selective) improves precision by leveraging generative models (e.g., ChatGPT) to identify continuous sensitive spans (e.g., PII), it remains limited to handling diverse unlearning targets (e.g., copyrighted content). This rigidity in identification compromises both effectiveness and generalizability. (C3) Sensitivity to Forget Set Size. Methods based on preference optimization (zhang2024negative; rafailov2024direct) mitigate catastrophic collapse more effectively than other baselines, helping to preserve model utility. However, their effectiveness declines significantly as the forget set size increases, resulting in notable utility loss (liu2024large), as shown in Figure 2.
Figure 2: Comparison of our TPO and NPO (zhang2024negative) on key metrics: forget quality and model utility. The results are derived from the Forget05 task of the TOFU dataset (maini2024tofu).
Figure 3: Illustration of responses generated by models unlearned using TPO (ours) and NPO on the retain set. The models were unlearned on the Forget05 task of the TOFU.
@@
To address the challenges (C1)-(C3), we propose the Targeted Information Forgetting (TIF) framework, as illustrated in Figure 1. Our main contributions are listed below:
@@
(1) TIF Framework. To tackle (C1), we propose a novel TIF framework for LLM unlearning. Different from existing unlearning approaches such as NPO, which predominantly unlearn entire information associated with the forget instances (e.g., $\bf{A_{f}}$ in the left panel of Figure 1), our TIF is designed to unlearn only the targeted unwanted information, such as the city of born in the right panel’s example. General information is often associated with some “General Words (GW)”, including stop words and commonly used phrases, which frequently appear in both retain and forget sets. In contrast, “Unwanted Words (UW)” correspond to specific private or copyrighted content, such as city of born. By specifically targeting only UW for unlearning, our TIF preserves more general information compared to existing methods like NPO, effectively preventing over-forgetting and enabling the model to generate more readable responses, as demonstrated in the retain set answers in Figure 1.
@@
(2) Unwanted Information Identification. To address (C2), we develop flexible yet effective approaches for unwanted information identification: a generative model such as ChatGPT-4, and a discriminative model such as DistilBERT (sanh2019distilbert), to effectively differentiate UW from GW. We evaluate their unlearning performance and illustrate their respective use cases. As a bottom line, even identifying function words (e.g., the, is, or an) as GW according to linguistics would improve model utility preservation.
@@
(3) A Novel Optimization Method to Retain Model Utility. To overcome (C3), we advance preference optimization algorithms  (zhang2024negative; fan2024simplicity) by introducing Targeted Preference Optimization (TPO), a novel optimization objective designed to mitigate the significant utility degradation observed in NPO. Specifically, our TPO integrates two innovative components: Preservation loss (PL) to maintain general model utility by retraining on GW, and Logit preference loss (LPL) to unlearn unwanted information in UW. This optimization approach effectively balances general information retention and unwanted information forgetting, improving the robustness of preference-guided optimization even with larger forget sets. As shown in Figure 2, our approach, TPO, achieves a comparable forget quality to NPO while significantly preserving a higher model utility. This allows the model to generate accurate information for answers in the retain set. In contrast, NPO struggles to retain essential knowledge from the retain set, as illustrated in Figure 3.
Problem Formulation
LLM Unlearning
@@
To achieve this goal, the unlearning procedure incorporates a forgetting objective on $\mathcal{D}_{f}$ and a retention objective on $\mathcal{D}_{r}$. Formally, the unlearning process is defined as (yao2023large; fan2024simplicity):
$\min_{\boldsymbol{\theta}}\mathbb{E}_{\xi_{f}\sim\mathcal{D}_{f}}[\ell_{f}(y_{f}|x_{f};\boldsymbol{\theta})]+\mathbb{E}_{\xi_{r}\sim\mathcal{D}_{r}}[\ell_{r}(y_{r}|x_{r};\boldsymbol{\theta})],$ (1)
where $\ell_{f}$ and $\ell_{r}$ represent the forget and retain losses, respectively. Specifically, the forget loss $\ell_{f}$ determines how well the model $\mathcal{M}_{\boldsymbol{\theta}}$ suppresses the association between $x_{f}$ and $y_{f}$, ensuring unwanted information is unlearned. Meanwhile, the retain loss $\ell_{r}$ enhances the model’s ability to maintain accurate associations between $x_{r}$ and $y_{r}$, preserving its original performance on $\mathcal{D}_{r}$.
Targeted Unlearning
As discussed earlier, a majority of works (si2023knowledge; chen2023unlearn; zhang2024negative; fan2024simplicity) have consistently treated the entire token sequence $y_{f}$ as the unlearning target for each sample $\xi_{f}\coloneqq(x_{f},y_{f})$ in the forget set $\mathcal{D}_{f}$, overlooking a critical question central to the process of LLM unlearning. Are all the words in the forget sample essential for unlearning in LLMs?
@@
$\min_{\boldsymbol{\theta}}\mathbb{E}_{\xi_{f}\sim\mathcal{D}_{f}}[\ell_{f}(\hat{y}|x_{f};\boldsymbol{\theta})]+\mathbb{E}_{\xi_{r}\sim\mathcal{D}_{r}}[\ell_{r}(y_{r}|x_{r};\boldsymbol{\theta})],$ (2)
@@
Targeted Information Forgetting (TIF)
@@
Unwanted Information Identification
@@
Discriminative Encoder-Only LM.
To detect unwanted information for unlearning tasks, we utilize an encoder-only LM, DistilBERT (sanh2019distilbert), denoted as $\mathcal{M}_{\text{bert}}$. This method leverages the contextual encoding of masked LMs to estimate the likelihood of each masked word, allowing differentiation between GW and UW. Firstly, given a sample $\xi_{f}\coloneqq(x_{f},y_{f})$ from the forget set, where $y_{f}=[w_{1},\cdots,w_{i},\cdots,w_{n}]$ is a word sequence, we sequentially replace each word $w_{i}$ in $y_{f}$ with a special [MASK] token. This transformation produces a masked sequence $y^{\prime}_{f_{i}}=[w_{1},\dots,w^{\prime}_{i},\dots,w_{n}]$, where $w^{\prime}_{i}=\texttt{[MASK]}$, as illustrated in the left panel of Figure 4. Next, the masked sequence $y^{\prime}_{f_{i}}$ is fed into $\mathcal{M}_{\text{bert}}$ along with $x_{f}$ to predict the masked token, formally: $w^{\text{pred}}_{i}=\mathcal{M}_{\text{bert}}(x_{f},y^{\prime}_{f_{i}})$. If the predicted masked token matches the original masked word, $w_{i}$ is labeled as GW, indicating general information. Otherwise, $w_{i}$ is marked as UW for target unlearning.
Generative Decoder-Only LM.
To harness the power of generative decoder-only LMs in capturing contextual and semantic information from text, we employ ChatGPT-4 to directly distinguish between UW and GW by analyzing the semantics of $y_{f}$, as shown in the right panel of Figure 4. Detailed task instructions can be found in Table 7 in the Appendix F. Furthermore, we also present a detailed comparison of discriminative and generative approaches in relation to unlearning performance in Appendix D.
Figure 5: Model utility across various methods on the TOFU forget05 task. Each line represents evaluations conducted at every epoch (6 steps). “–GPT” denotes the use of ChatGPT-4 for unwanted information identification, while “PL” refers to the approach plus the PL Loss.
Figure 6: Forget quality across various methods on the TOFU forget05 task. Each line represents evaluations conducted at every epoch (6 steps). “–GPT” denotes the use of ChatGPT-4 for unwanted information identification, while “PL” refers to the approach plus the PL Loss.
Targeted Preference Optimization (TPO)
Motivation. Although numerous unlearning methods, such as NPO (zhang2024negative), have demonstrated strong performance on benchmarks such as TOFU (maini2024tofu) and MUSE (shi2024muse), most of them struggle with model utility degradation when handling large forget sets (liu2024large). Specifically, we evaluate the model utility and forgetting quality of NPO on TOFU, as illustrated in Figures 5 and 6. A significant decline in model utility is evident, with the score dropping from 0.76 to 0.11, highlighting a severe utility degradation issue. We hypothesize that this degradation stems from NPO’s indiscriminate handling of the entire forget samples as unlearning targets, failing to differentiate between unwanted and general information.
@@
Preservation Loss (PL). To further investigate the reason for the model utility degradation observed in NPO-GPT, we analyze the cross-entropy loss values for GW, while the unlearning is limited to UW, as defined in Equation (2). Although the model is not explicitly optimized to forget GW, the increase in loss values for NPO-GPT indicates that GW are also forgotten, as shown in Figure 7. This observation naturally leads to our key idea: how can we prevent the forgetting effects on GW and, in turn, keep their loss values as low as possible during unlearning optimization? To address this, we introduce the Preservation Loss (PL), which integrates a cross-entropy loss term on GW to explicitly prevent the model from forgetting general information, formally:
@@
To validate the effectiveness of the PL term, we integrate it into NPO-GPT, forming NPO+PL-GPT, and evaluate its performance. As illustrated in Figure 7, incorporating PL helps maintain stable and low loss values for GW. Consequently, NPO+PL-GPT exhibits a significantly slower decline in model utility while achieving comparable forget quality to NPO and NPO-GPT, as shown in Figures 5 and 6. These initial results demonstrate that PL effectively mitigates model utility degradation, particularly for preserving the general information we aim to retain.
Figure 7: Evaluation of cross-entropy loss values on GW for NPO-GPT and NPO-GPT+PL at each unlearning step. All results are obtained for the Forget05 task in the TOFU dataset, with models trained over 10 epochs.
Figure 8: The KL divergence between the reference model and the unlearned models on both forget and retain sets. All results are obtained for the Forget05 task in the TOFU dataset, with models trained over 10 epochs.
Logit Preference Loss (LPL). Upon further examination of Figure 5, we observe that model utility degradation persists even after incorporating unwanted information identification (GPT) and PL into NPO, as seen in NPO+PL-GPT. We hypothesize that this issue stems from the limitations of NPO itself. Specifically, the unlearning process in NPO likely introduces excessive changes from the original model $\mathcal{M}_{\boldsymbol{\theta}_{\text{o}}}$, which serves as a reference model with parameters kept frozen during unlearning, to the final unlearned model $\mathcal{M}_{\boldsymbol{\theta}}$, particularly affecting certain general information.
To further validate this hypothesis, we analyze the logit distribution by computing the KL divergence between $\mathcal{M}_{\boldsymbol{\theta}_{\text{o}}}$ and $\mathcal{M}_{\boldsymbol{\theta}}$ for both the forget and retain sets, as shown in Figure 8. The high KL divergence observed in both sets suggests that $\mathcal{M}_{\boldsymbol{\theta}}$ (NPO-GPT+PL) has unintentionally forgotten not only the unwanted information but also general information. Specifically, NPO functions by directly reducing the probability assigned to target tokens, which is computed using the softmax function: $P(y_{t})=\frac{\exp{(z_{t})}}{\sum^{V}_{j=1}\exp(z_{j})}$, where the $V$ represents the vocabulary size, and $z_{t}$ denotes the logit for target token $y_{t}$. However, reducing the probability of target token $P(y_{t})$ can be achieved not only by decreasing its logits $z_{t}$ but also by increasing the logits of other tokens in the vocabulary. This unintended effect distorts the model’s overall logit distribution, potentially compromising its ability to retain general information. The key challenge is to develop a new optimization strategy for the target tokens (UW) that selectively impacts their logit distribution while preserving the general information from GW.
@@
$\ell_{\text{LPL}}(\boldsymbol{\theta})=-\mathbb{E}_{\xi_{f}\sim D_{f}}\bigg[\frac{2}{\beta}\log\sigma\bigg(\beta\frac{1}{|\hat{y}|}\sum_{i=1}^{|\hat{y}|}(z_{i}^{\boldsymbol{\theta}_{\text{o}}}-z_{i}^{\boldsymbol{\theta}})\bigg)\bigg],$ (4)
where $z_{i}$ here denotes the logit of target token $\hat{y}_{i}$, and $\boldsymbol{\theta}$ and $\boldsymbol{\theta}_{\text{o}}$ represent the parameters of the unlearned model $\mathcal{M}_{\boldsymbol{\theta}}$ and the original model $\mathcal{M}_{\boldsymbol{\theta}_{\text{o}}}$, respectively.
Different from NPO, LPL explicitly reduces only the logits of target tokens (UW) by enforcing a preference loss between $\mathcal{M}_{\boldsymbol{\theta}}$ and $\mathcal{M}_{\boldsymbol{\theta}_{\text{o}}}$. The primary function of LPL is to maximize the difference in target token logits between $\mathcal{M}_{\boldsymbol{\theta}}$ and $\mathcal{M}_{\boldsymbol{\theta}_{\text{o}}}$, while preserving the overall logit distribution. This targeted approach ensures that only the unwanted information associated with the target tokens is unlearned, without affecting general information across other tokens. As a result, LPL enables a more precise unlearning process while significantly improving model utility retention.
@@
$\begin{split}\mathbb{E}_{\xi_{f}\sim D_{f}}\bigg[&-\tfrac{2}{\beta}\log\sigma\big(\beta\,(z_{\boldsymbol{\theta}_{\text{o}}}(\hat{y}|x_{f})-z_{\boldsymbol{\theta}}(\hat{y}|x_{f}))\big)\\ &-\log P_{\boldsymbol{\theta}}(\bar{y}\mid x_{f})\bigg],\end{split}$ (5)
@@
Finally, as the initial results shown in Figure 8, our approach TPO-GPT minimizes the disruption to the logit distribution on both the forget and retain sets compared to NPO+PL-GPT. Furthermore, TPO-GPT maintains most of the model utility while achieving a comparable level of forget quality to the NPO-based methods, as shown in Figures 5 and 6.
Experimental Setting
Datasets and Metrics
@@
We evaluate the proposed approach alongside the baseline methods on the two widely used benchmark datasets: MUSE (shi2024muse) and TOFU (maini2024tofu).